OpenStudy (anonymous):

A wildlife biologist caught, banded, and released 300 quail at a local quail ranch. One month later, he captured 100 quail in the same reserve and discovered that 14 of these were banded. Estimate the number of quail on the quail ranch and match your result to the correct answer below. Select one: a. 1700 b. 1400 c. 600 d. 4200

OpenStudy (campbell_st):

well this is capture recapture that uses the ratio of banded against the population . let p be the population then \[\frac{p}{300} = \frac{100}{14}\] just solve for p

OpenStudy (anonymous):

2142.85 @campbell_st

OpenStudy (campbell_st):

that's what i got... and its interesting that you don't have an answer choice matching that

OpenStudy (campbell_st):

can you just check the information is correct... and the answer choices are correct..

OpenStudy (anonymous):

they are correct i just check @campbell_st

OpenStudy (campbell_st):

ummmm well I now have no clue... the method is correct...

OpenStudy (anonymous):

Christa bought a total of 6 lb of flounder and shrimp and paid $18 for the flounder and $18 for the shrimp. If the price per pound of the shrimp is twice that of the flounder, how many pounds of each did she buy? Select one: a. 4 lb shrimp and 2 lb flounder b. 2 lb shrimp and 4 lb flounder c. 3 lb shrimp and 3 lb flounder d. 1 lb shrimp and 5 lb flounder how about this one than @campbell_st

OpenStudy (campbell_st):

4lb flounder cost per pound = 18/4 = $4.50 2lb shrimp cost per pound = 18/2 = $9.00

OpenStudy (anonymous):

you are a freaking lifesaver David hires a septic system contractor to dig trenches for his new septic tank. The cost to hire the contractor is a fixed cost of $130 plus an additional $85 per hour. Which expression below describes the average cost of an x-hour job? @campbell_st

OpenStudy (campbell_st):

C = 85t + 130 t = number of hours C = cost

OpenStudy (anonymous):

thanks im still stuck on the wildfire one ugh @campbell_st

OpenStudy (campbell_st):

well I'd say the creator of the question has made a mistake.... happens a fair bit...

OpenStudy (campbell_st):

if there 200 tagged quail it comes down to 1400